# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v2

@@ C.2 Dominant Sector Comparison Between the Strawman Approach and TraCS-D (Example 1)
Figure 12: Comparison of the size and probability of the dominant sector between TraCS-D and the strawman approach. TraCS-D (solid line) achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach (dashed line), having better adaptiveness across different $\varepsilon$ regions.
@@
C.3 Detailed Form of $|\mathcal{D}_{r(\varphi)}|$ (Section 3.2.2)
@@
C.4 More Extensions (Section 3.5.5)
C.4.1 Other Piecewise-based Mechanisms
In Definition 5 and Definition 6, we use specific instantiations for the parameters $p$ and $[l,r)$ in the piecewise-based mechanism. Besides these instantiations, TraCS can be integrated with any other piecewise-based mechanisms [31, 21, 24, 23, 38] by redesigning them. For example, SW [21] is a famous mechanism proposed for distribution estimation; it is defined on $0,1)\to[-b,1+b)$, with $b$ as a parameter. It uses different $p$ and $[l,r)$, and employs a strategy of maximizing the mutual information between the perturbed and sensitive data to determine $b$. We can redesign SW for the direction perturbation and distance perturbation in TraCS-D and TraCS-C. Appendix C.5 provides the detailed redesign of SW for them and shows intuitive comparisons with Definition 5 and Definition 6. The experimental section also provides an comparison between the redesigned SW and TraCS-D.
C.4.2 Other Shapes of Location Space
The design of LDP mechanisms is space-specific, implying that extensions to other shapes of location space are non-trivial.
For TraCS, the challenge lies in determining the size of each subspace, i.e. $|\mathcal{D}_{\varphi}|$, $|\mathcal{D}_{r(\varphi)}|$, $|\mathcal{D}_{a}|$, and $|\mathcal{D}_{b}|$. Among them, $|\mathcal{D}_{r(\varphi)}|$ is not apparent but can be calculated for rectangular location spaces. Although calculating $|\mathcal{D}_{r(\varphi)}|$ is generally non-trivial for irregular shapes, TraCS can be extended to parallelogram location spaces by a linear transformation to a rectangular space. Meanwhile, TraCS-D can be used for circular areas, which are widely used in relative location representation, e.g. radar and sonar systems.
TraCS-D for circular area. When the location space is circular with radius $R$, TraCS-D has a more straightforward design. The main insight is that such a location space $\mathcal{S}_{R}$ has a natural representation in polar coordinates, i.e. $\mathcal{S}_{R}=\{(\varphi,r):\varphi\in[0,2\pi),r\in[0,R)\}$. Any location $\tau_{i}\in\mathcal{S}_{R}$ corresponds to a pair of coordinates $(\varphi_{i},r_{i})$, and this representation is unique and reversible. In this case, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ and the distance space $\mathcal{D}_{r}=[0,R)$ are independent, so we can apply $\mathcal{M}_{\circ}$ to $\varphi_{i}$ and $\mathcal{M}_{-}$ to $r_{i}/R$. The perturbed location $\tau_{i}^{\prime}$ is then $(\varphi_{i}^{\prime},r_{i}^{\prime})=(\mathcal{M}_{\circ}(\varphi_{i}),\mathcal{M}_{-}(r_{i}/R)\cdot R)$, still within polar coordinates.
C.5 Redesigned SW (Appendix C.4.1)
@@ C.4.2 Other Shapes of Location Space
C.6 Experimental Results for Extensions (Appendix C.4)
C.6.1 Effect of Different Piecewise-based Mechanisms
@@
C.6.2 TraCS-D for Circular Area
@@
Figure 13: Comparison between different piecewise-based mechanisms.
Figure 14: Fifty samples of TraCS-D at $\tau=(\pi,0.5)$ for circular area.
C.7 Experimental Results for Other Metrics (Section 4.1.1)
A smaller $AE$ generally indicates better preservation of range queries and hotspots [9, 37]. (i) For range query preservation (Formula 17 in [9]), given a threshold $\delta$, a perturbed location is considered correct if it lies within a $\delta$ distance of the corresponding sensitive location. Thus, when $AE\leq\delta$, the perturbed location is expected to satisfy the range query, and a smaller $AE$ implies higher correctness. (ii) For hotspot preservation (Section 6.2.4 in [37]), which measures how many hotspots from a given set remain after perturbation, a smaller $AE$ leads to better hotspot retention after rounding in TraCS.
(a) Range query preservation
(b) Hotspot preservation
Figure 15: Comparison of other trajectory metrics on the CHI dataset. Higher range query preservation and lower error in hotspot preservation indicate better performance.
  • Range query preservation: We set the threshold $\delta=0.1$, meaning a perturbed location is considered correct if its distance from the sensitive location is less than $0.1$ in longitude-latitude coordinates.
  • Hotspot preservation: The hotspot set is defined as the first $50\%$ of locations in the CHI dataset. We count the number of hotspots before and after perturbation for the first $100$ trajectories and report the difference.
The results are shown in Figure 15. It can be seen that TraCS generally outperforms existing methods across different privacy levels, demonstrating superior effectiveness in preserving both range queries and hotspots. Particularly, TraCS shows significant improvement in hotspot preservation compared to existing methods across all privacy levels.

### v3

@@ C.4. Dominant Sector Comparison Between the Strawman Approach and TraCS-D (Example 2)
Figure 14. Comparison of the size and probability of the dominant sector between TraCS-D and the strawman approach. TraCS-D (solid line) achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach (dashed line), having better adaptiveness across different $\varepsilon$ regions.
@@
C.5. Detailed Form of $|\mathcal{D}_{r(\varphi)}|$ (Section 3.2.2)
@@
C.6. Redesigned SW (Section 3.5.7)
@@
C.7. Experimental Results for Extensions (Section 3.5.7)
C.7.1. Effect of Different Piecewise-based Mechanisms
@@
C.7.2. TraCS-D for Circular Area
@@
Figure 15. Different piecewise-based mechanisms.
Figure 16. Samples of TraCS-D at $\tau=(\pi,0.5)$ for circular area.
C.8. 2D Laplace Mechanism with Truncation (Section 4.1.1)
For 2D location spaces equipped with the Euclidean distance, a standard Laplace-based approach is the Planar Laplace mechanism (DBLP:conf/ccs/AndresBCP13). [^2] We omit the standard (truncated) Laplace mechanism for $\mathbb{R}$, as it is designed for $L_{1}$ distance and has been shown to have worse data utility compared to piecewise-based mechanisms (DBLP:journals/popets/ZhengMH25). It was originally proposed for geo-indistinguishability, and it also provides an LDP guarantee with a calibrated sensitivity.
Definition 0 (Planar Laplace mechanism).
Let $\mathcal{X}\subseteq\mathbb{R}^{2}$ be the input domain. The Planar Laplace mechanism $\mathcal{M}_{\mathrm{PL}}:\mathcal{X}\to\mathbb{R}^{2}$ is defined by
$\mathcal{M}_{\mathrm{PL}}(x)=x+\eta,$
where $\eta\in\mathbb{R}^{2}$ is a noise vector with density
$pdf[\eta]=\frac{\varepsilon^{2}}{2\pi}\exp\bigl(-\varepsilon\|\eta\|_{2}\bigr).$
Sensitivity calibration for an LDP guarantee. By the triangle inequality, for any $x_{1},x_{2}\in\mathcal{X}$ and any $y\in\mathbb{R}^{2}$, we have
$\frac{pdf[y-x_{1}]}{pdf[y-x_{2}]}=\exp\Bigl(\varepsilon\bigl(\|y-x_{2}\|_{2}-\|y-x_{1}\|_{2}\bigr)\Bigr)\leq\exp\bigl(\varepsilon\|x_{1}-x_{2}\|_{2}\bigr).$
Therefore, over the domain $\mathcal{X}$, the mechanism satisfies $(\varepsilon\cdot\mathrm{diam}(\mathcal{X}))$-LDP, where $\mathrm{diam}(\mathcal{X})$ is the diameter of $\mathcal{X}$. Equivalently, to ensure $\varepsilon$-LDP for all pairs of inputs in $\mathcal{X}$, one can run the Planar Laplace mechanism with parameter $\varepsilon/\mathrm{diam}(\mathcal{X})$.
Truncation for bounded location spaces. The Planar Laplace mechanism outputs perturbed locations in $\mathbb{R}^{2}$, which may fall outside the bounded location space $\mathcal{S}$. A common practice is to truncate the output to $\mathcal{S}$, e.g. by projecting it to the nearest point in $\mathcal{S}$. This truncation is a post-processing step and does not affect the LDP guarantee.
C.9. Privacy Parameter Assignment in TraCS-D (Section 4.1.1)
In TraCS-D, the overall perturbation error is jointly determined by the direction perturbation and the distance perturbation. These two components are controlled by $\varepsilon_{d}$ for $\mathcal{M}_{\circ}$ and $\varepsilon-\varepsilon_{d}$ for $\mathcal{M}_{-}$, respectively. As a result, the (theoretical) error bound of TraCS-D depends on both $\varepsilon$ and the privacy split $\varepsilon_{d}$. In principle, the best choice of $\varepsilon_{d}$ is the one that minimizes this bound.
However, the optimal split depends on $\varepsilon$ and on the shape of $\mathcal{S}$, which makes a universal closed-form choice impractical. Using the same experimental setup as in the Figure 4(a), we empirically evaluate how $\varepsilon_{d}$ affects the error of TraCS-D; the results are shown in the figure on the right. We observe a consistent pattern: for each fixed $\varepsilon$, the error decreases initially and then increases as $\varepsilon_{d}$ varies from 0 to the full budget $\varepsilon$.
Across all tested $\varepsilon$ values, the minimizer $\varepsilon_{d}/\varepsilon$ is close to our heuristic split $\varepsilon_{d}=\pi/(1+\pi)$ (dashed pink line), which supports the use of this heuristic in practice.
C.10. Privacy Parameter Assignment for a Whole Trajectory (Section 4.2.6)
(a) TKY dataset
(b) CHI dataset
Figure 17. Comparison with trajectory-level $\varepsilon$ assignment. The average trajectory length is $113$ for TKY and $13$ for CHI. TraCS’s AE decreases fast as $\varepsilon$ increases, outperforming other discrete-space mechanisms when $\varepsilon$ is large.
(a) TKY location space
(b) CHI location space
Figure 18. Location spaces for TKY and CHI datasets. TraCS operates over the whole rectangular area that encloses the discrete location space of each city, while other discrete-space mechanisms operate only on the discrete locations (blue points).
We compare TraCS with NGram, L-SRR, and ATP under a trajectory-level privacy parameter assignment on the TKY and CHI datasets. The average trajectory lengths are $113$ (TKY) and $13$ (CHI). Accordingly, we assign a trajectory-level budget of $\varepsilon\times 113$ for TKY and $\varepsilon\times 13$ for CHI. This scaling makes the trajectory-level setting more comparable to the location-level $\varepsilon$ assignment in Figure 9. The results are shown in Figure 17.
We observe that TraCS has larger AEs than the discrete-space mechanisms when $\varepsilon$ is small, but its errors decrease rapidly as $\varepsilon$ increases. As a result, TraCS outperforms the discrete-space mechanisms in the large-$\varepsilon$ regime. Compared with assigning $\varepsilon$ at the location level in Figure 9, the overall trends remain similar across all mechanisms: TraCS starts to outperform all other discrete-space mechanisms when $\varepsilon\approx 4$ per location on both datasets. Among the discrete-space mechanisms, ATP achieves the lowest AEs on TKY, whereas NGram (i.e. the Exponential mechanism with a reachable set) attains the lowest AEs on CHI, due to CHI’s smaller location space than TKY.
As discussed in Section 4.2.3, the larger average errors (AEs) of TraCS in the small-$\varepsilon$ regime can be attributed to its enlarged effective location space. TraCS is designed for continuous spaces and perturbs locations over the rectangular area that encloses the discrete location space of each city, as illustrated in Figure 18. The discrete location set in TKY is comparatively dense and evenly distributed, whereas the location set in CHI is much sparser. Consequently, the gap between the discrete set and its enclosing rectangle is larger for CHI, which amplifies the utility loss of TraCS and explains its weaker performance on CHI (relative to discrete-space mechanisms) compared with TKY.